OpenStudy (anonymous):

sample of Geometry class final exam grades are as follows: 80, 85, 72, 76, 95 Find the mean absolute deviation

OpenStudy (dumbcow):

do you mean standard deviation?

OpenStudy (dumbcow):

\[mean = \frac{80+85+72+76+95}{5} = 81.6\] \[stdev = \sqrt{\frac{\sum_{}^{}(x _{i}-mean)^{2}}{n-1}} = 8.905\]

OpenStudy (anonymous):

mean=81.6 so mean absolute deviation is \[\frac {|80-81.6|+|85-81.6|+|72-81.6|+|76-81.6|+|95-81.6|}{5}\]